OpenStudy (anonymous):

Line AB contains points A (−2, 3) and B (4, 5). Line AB has a slope that is zero undefined positive negative

OpenStudy (anonymous):

use the slope formula slope = (y2 - y1) / (x2 - x1) (-2,3) x1 = -2 and y1 = 3 (4,5) x2 = 4 and y2 = 5 now we sub slope = (5 - 3) / (4 - (-2) now do the math and find the slope :)